# Asymmetric Contrastive Multimodal Learning for Advancing Chemical Understanding

### v3

@@ 2.5 Molecular property prediction tasks
Molecular property prediction is a critical graph-level prediction task used to assess the generalization capability of pre-trained graph encoders. In other words, it evaluates how well the encoder’s learned representations can be adapted to downstream tasks. In this section, we utilize the pretrained graph encoder from the ACML framework for downstream fine-tuning tasks, incorporating 6 classification and 2 regression benchmarks using MoleculeNet datasets [50]. For each task, we employ the shallow GIN [51] with 5 layers and 64 hidden dimensions as the backbone graph encoder, which is initialized with weights pre-trained by the ACML framework. We compare our models with other pretraining baselines on molecular graph learning in both classification and regression tasks. For classification tasks, ROC-AUC is used as the evaluation metric, and for regression tasks, RMSE is used. Among the baselines, L2P-GNN [52] and MGSSL [53] with different generation orders (BFS and DFS), MotifConv [44] propose self-supervised tasks within graphs to capture sub-structural or motif-level information during pretraining, while MolCLR [30] and MoMu [54] develop contrastive learning using augmented graphs or other modalities during pretraining. For a fair comparison, all baselines also use the 5-layer GIN model as the backbone graph encoder. We use scaffold splitting [55] to divide the dataset into the training set, validation set, and test set with a ratio of 8:1:1. For each dataset, the average and standard deviation of 5 individual runs are reported. More experimental details can be found in Section 5.2.2.
@@
Table 2 shows the test performance on 8 molecular benchmarks. First, by comparing with no pretrained models, we observed that the best ACMLs generally outperform models without pretraining by big margins. Specifically, G-Image makes improvements in all datasets, including an average increase in AUC-ROC of 10.0% in BACE, 16.4% in Clintox, 7.9% in HIV, and reductions in RMSE with a 33.3% average decrease in ESOL and a 47.9% decrease in FreeSolv. These results suggest that contrastive learning with G-Image captures useful information in molecular depictions for a better understanding of the overall molecular structure. Second, by comparing with other self-supervised pretraining strategies, the proposed ACML framework achieves the best performance on 7 of 8 datasets. These improvements indicate that our ACML is a powerful self-supervised learning strategy, which is easy to implement and requires minimal manual design of self-supervised tasks. Third, by comparing ACMLs using different chemical modalities, we observed that G-Image generally outperforms other modalities as it achieves the best performance in BACE, BBBP, Sider, and ESOL, and ranks second in FreeSolv. In the Tox21 and HIV datasets, G-Image’s performance is not the highest, but it closely approaches the top results. This is likely due to its ability to capture comprehensive molecular structure information, which is crucial for most molecular property prediction tasks. These observations on the superiority of G-Image are also consistent with the analysis in Section 2.4: graph representations from G-Image exhibit high correlations with various chemical properties in drug discovery. In contrast, other modalities provide specific insights that could be beneficial for certain tasks but not universally effective. For instance, G-1H NMR achieves the highest average ROC-AUC of 83.9% in Clintox, 81.6% in Tox21, and 84.9% in HIV, but performs worse compared to the no-pratrained models in BBBP (3.3%) and Sider (3.3%). Similar cases are observed for other tasks and other modalities.
@@
BACE $\uparrow$ BBBP $\uparrow$ Clinton $\uparrow$ Sider $\uparrow$ Tox21 $\uparrow$ HIV $\uparrow$ Esol $\downarrow$ FreeSolv $\downarrow$
GIN (no pretrain) [51] 79.7 $\pm$ 4.9 87.3 $\pm$ 3.6 53.0 $\pm$ 6.5 61.6 $\pm$ 2.5 78.3 $\pm$ 2.4 76.2 $\pm$ 5.8 1.262 $\pm$ 0.156 4.672 $\pm$ 1.549
L2P-GNN [52] 81.9 $\pm$ 0.4 87.0 $\pm$ 0.8 54.0 $\pm$ 2.4 61.7 $\pm$ 1.8 77.4 $\pm$ 0.6 78.0 $\pm$ 2.6 1.772 $\pm$ 0.479 3.733 $\pm$ 1.323
MGSSL (DFS) [53] 79.7 $\pm$ 0.8 70.5 $\pm$ 1.1 79.7 $\pm$ 2.2 60.5 $\pm$ 0.7 76.4 $\pm$ 0.4 79.5 $\pm$ 1.1 0.914 $\pm$ 0.095 3.401 $\pm$ 0.896
MGSSL (BFS) [53] 79.1 $\pm$ 0.9 69.7 $\pm$ 0.1 80.7 $\pm$ 2.1 61.8 $\pm$ 0.8 76.5 $\pm$ 0.3 78.8 $\pm$ 1.2 0.861 $\pm$ 0.056 2.596 $\pm$ 0.997
MotifConv [44] 82.0 $\pm$ 5.5 90.0 $\pm$ 3.1 65.5 $\pm$ 13.9 62.7 $\pm$ 2.8 80.2 $\pm$ 1.5 80.0 $\pm$ 4.3 1.401 $\pm$ 0.153 3.697 $\pm$ 0.400
MoMu [54] 72.1 $\pm$ 0.3 70.2 $\pm$ 0.4 71.1 $\pm$ 3.0 56.1 $\pm$ 0.4 74.5 $\pm$ 0.5 77.2 $\pm$ 1.5 1.413 $\pm$ 0.104 3.374 $\pm$ 0.752
MolCLR [30] 74.6 $\pm$ 1.1 68.0 $\pm$ 2.0 62.0 $\pm$ 2.7 56.3 $\pm$ 2.2 73.9 $\pm$ 0.5 72.1 $\pm$ 1.3 1.354 $\pm$ 0.213 2.817 $\pm$ 0.911
@@ 3 Discussion and Conclusion
In this work, we introduced ACML framework, a novel approach tailored for molecular representational learning based on multi-modality contrastive learning. We demonstrated the inheritance of molecular graph-like structure allows graph representation to convey extensive information about molecules and treat the graph neural networks, i.e., the graph encoder, as a receptor to assimilate chemical semantics through multi-modal coordination, leading to effective and explainable molecular representational learning. Extensive experimental results on important chemical tasks, such as isomer discrimination, uncovering crucial chemical properties for drug discovery, and molecular properties prediction from MoleculeNet datasets, demonstrate the cross-modality transfer ability and the power of graph neural networks. In addition, our findings reveal a series of interesting questions to consider: of particular interest is how to effectively train a graph encoder to make it tailored to molecular representational learning. Instead of designing a deep and complex GNN framework and proposing various levels of self-supervised pretraining tasks, we demonstrate that the use of multi-modality contrastive learning can allow a shallow GNN with light training resources to produce an expressive graph encoder with great interpretability.
4 Methods
4.1 The ACML framework
@@
First, the inputs of each ACML framework consist of a molecular graph modality and one chemical modality. For each chemical modality, we make use of effective pre-trained encoders from publicly available sources, which have been demonstrated to be effective in the corresponding downstream tasks (See Table 3). As for molecular graph representation, we employ the graph neural network (GNN) as the graph encoder, where we elaborated several GNN variants in Section 4.2. These inputs are then fed into their respective encoders to produce embeddings. Mathematically, let $g$ and $c$ denote a molecule’s graph representation and its chemical modality representation, their embeddings after passing to their corresponding encoders are:
@@
Table 3: Unimodal encoders of pervasive chemical modalities.
4.2 Graph encoder
Chemical molecules can be naturally represented as attributed relational graphs $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are the node set and the edge set. Here a node $v\in\mathcal{V}$ and an edge $(v,u)\in\mathcal{E}$ represent the atom $v$ and the chemical bond connecting $u$ and $v$, respectively. The corresponding atom features (atomic number, chirality tag, hybridization, etc,.) and bond features (bond types, stereotypes, etc,.) are treated as node attributes and edge attributes in the graph. Graph neural networks (GNNs) [58, 59] which operate on graphs, have been combined with deep learning [60] to learn representations on graph data [61, 62, 63]. GNNs are also demonstrated as effective frameworks for molecular representational learning [64, 39].
Graph convolution. Many popular graph convolutional frameworks follow the message-passing scheme, including an iterative way of updating node representations based on the aggregations from neighboring nodes. Widely used approaches, such as GCN [65], GIN [51], GAT [66] and GraphSage [67], all follow message passing schemes. For the rest of the paper, we refer “message passing based GNNs” as “GNN”. For simplicity, we denote $\mathbf{x}_{u}^{(k)}$ as node feature vector of node $u$ in $k$-th layer, $\mathbf{e}_{v,u}$ as (optional) edge feature vector from node $v$ to node $u$, and let $\mathcal{N}_{u}$ represent the neighboring node set of the central node $u$. One message passing based convolution layer includes two steps: (1) create the message between the central node $u$ and its neighbor $v$, formalized by the function $\phi^{(k)}(\mathbf{x}_{u}^{(k)},\mathbf{x}_{v}^{(k)},\mathbf{e}_{v,u})$; (2) aggregate the messages of all neighbors of the central node $u$ using the operator $\bigoplus$, which denotes a differentiable and permutation invariant function, e.g., sum, mean or max operator. Mathematically, the latent node presentation is updated according to Equation 3 and will be fed into the next convolutional layer.

### v4

@@ 2.5 Molecular Property Prediction Tasks
Molecular property prediction is a critical graph-level prediction task used to assess the generalization capability of pre-trained graph encoders. In other words, it evaluates how well the encoder’s learned representations can be adapted to downstream tasks. Additionally, finetuning on a small dataset is challenging because models primarily depend on their pre-training capabilities in this scenario. By using a limited dataset, we can better assess the model’s representation learning capabilities and understand how well the pre-trained knowledge transfers to specific tasks. Thus, we intentionally choose to finetune the model using relatively small datasets to evaluate the pretrained graph encoder from the ACML framework across a range of downstream tasks.
2.5.1 Experimental Settings
Datasets.
We employ two key molecular benchmark resources. The first, MoleculeNet [54], is a commonly used benchmark specially designed for testing machine learning methods of molecular properties. The second, Therapeutics Data Commons (TDC) [55], is a recent platform providing molecular datasets and learning tasks for therapeutics. The molecular data of TDC are in SMILES format, we convert them into molecular graphs using OGB package [56]. Our focus is on tasks involving small molecule property prediction, with an emphasis on properties relevant to drug discovery. Therefore, in MoleculeNet, we consider 6 classification and 2 regression datasets. In TDC, we consider 12 binary classification datasets about (ADME) properties (absorption, distribution, metabolism, and excretion). A small-model drug needs to travel from the site of administration (e.g., oral) to the site of action (e.g., a tissue) and then decompose, and exit the body. To do that safely and efficaciously, the chemical is required to have numerous ideal ADME properties. For classification tasks, ROC-AUC is used as the evaluation metric, and for regression tasks, RMSE is used.
Models and Baselines.
@@ 2.5.1 Experimental Settings
For each task, we employ the GIN [57] with 5 layers as the backbone graph encoder, initialized with weights pre-trained by the ACML framework. We compare our models with other pretraining baselines on molecular graph learning. Among the baselines, L2P-GNN [51] and MGSSL [52] with different generation orders (BFS and DFS), MotifConv [44] propose self-supervised tasks within graphs to capture substructural or motif-level information during pretraining, while MolCLR [30] and MoMu [53] develop contrastive learning using augmented graphs or other modalities during pretraining. MotifConv was excluded from TDC experiments due to its high complexity during pretraining. For a fair comparison, all baselines also use the 5-layer GIN model as the backbone graph encoder. Experiments using GIN without pretraining are also conducted for comparison. More experimental details can be found in Section B.2.2. For each dataset, the average and standard deviation of model performance are reported based on 5 independent runs.
2.5.2 Results and Analysis
MoleculeNet Datasets.
@@ 2.5.2 Results and Analysis
Table 4 shows the test performance on MoleculeNet. First, by comparing models without pretrainnig, we observed that the best ACMLs generally outperform by big margins. Specifically, G-Image makes improvements in all datasets, including an average increase in AUC-ROC of 10.0% in BACE, 16.4% in Clintox, 7.9% in HIV, and reductions in RMSE with a 33.3% average decrease in ESOL and a 47.9% decrease in FreeSolv. These results suggest that contrastive learning with G-Image effectively captures useful information in molecular depictions, leading to a better understanding of the overall molecular structure. Second, when compared with other self-supervised pretraining strategies, the proposed ACML framework achieves the best performance on 7 of 8 datasets. These improvements indicate that our ACML framework is a powerful self-supervised learning strategy, which is easy to implement and requires minimal manual design of self-supervised tasks. Third, by comparing ACMLs using different chemical modalities as inputs, we observed that G-Image generally outperforms other modalities as it achieves the best performance in BACE, Sider, and ESOL, and ranks second in FreeSolv. Although G-Image does not lead in BBBP, Tox21 and HIV, its performance is close to the top. This is likely due to its ability to capture comprehensive molecular structure information, which is crucial for most molecular property prediction tasks. These observations on the superiority of G-Image are also consistent with the analysis in Section 2.4, where graph representations from G-Image exhibit strong correlations with various chemical properties relevant to drug discovery. In contrast, other modalities provide specific insights that could be beneficial for certain tasks but not universally effective. For instance, G-1H NMR achieves the highest average ROC-AUC of 83.9% in Clintox, 81.6% in Tox21, and 84.9% in HIV, but performs worse compared to the no-pre-trained models in BBBP (-3.3%) and Sider (-3.3%). Similar cases are observed for other tasks and other modalities.
@@
BACE $\uparrow$ BBBP $\uparrow$ Clintox $\uparrow$ Sider $\uparrow$ Tox21 $\uparrow$ HIV $\uparrow$ Esol $\downarrow$ FreeSolv $\downarrow$
GIN (no pretrain) 79.7 $\pm$ 4.9 87.3 $\pm$ 3.6 53.0 $\pm$ 6.5 61.6 $\pm$ 2.5 78.3 $\pm$ 2.4 76.2 $\pm$ 5.8 1.262 $\pm$ 0.156 4.672 $\pm$ 1.549
L2P-GNN 81.9 $\pm$ 0.4 87.0 $\pm$ 0.8 54.0 $\pm$ 2.4 61.7 $\pm$ 1.8 77.4 $\pm$ 0.6 78.0 $\pm$ 2.6 1.772 $\pm$ 0.479 3.733 $\pm$ 1.323
MGSSL (DFS) 79.7 $\pm$ 0.8 70.5 $\pm$ 1.1 79.7 $\pm$ 2.2 60.5 $\pm$ 0.7 76.4 $\pm$ 0.4 79.5 $\pm$ 1.1 0.914 $\pm$ 0.095 3.401 $\pm$ 0.896
MGSSL (BFS) 79.1 $\pm$ 0.9 69.7 $\pm$ 0.1 80.7 $\pm$ 2.1 61.8 $\pm$ 0.8 76.5 $\pm$ 0.3 78.8 $\pm$ 1.2 0.861 $\pm$ 0.056 2.596 $\pm$ 0.997
MotifConv 82.0 $\pm$ 5.5 90.0 $\pm$ 3.1 65.5 $\pm$ 13.9 62.7 $\pm$ 2.8 80.2 $\pm$ 1.5 80.0 $\pm$ 4.3 1.401 $\pm$ 0.153 3.697 $\pm$ 0.400
MoMu 72.1 $\pm$ 0.3 70.2 $\pm$ 0.4 71.1 $\pm$ 3.0 56.1 $\pm$ 0.4 74.5 $\pm$ 0.5 77.2 $\pm$ 1.5 1.413 $\pm$ 0.104 3.374 $\pm$ 0.752
MolCLR 74.6 $\pm$ 1.1 68.0 $\pm$ 2.0 62.0 $\pm$ 2.7 56.3 $\pm$ 2.2 73.9 $\pm$ 0.5 72.1 $\pm$ 1.3 1.354 $\pm$ 0.213 2.817 $\pm$ 0.911
@@
Table 5: Compare test performance (mean $\pm$ std) on ADME properties prediction benchmarks of TDC. ROC-AUC (%) is used for evaluations. The best result for each dataset is in bold, with the second best underlined.
HIA PAMPA Permeability Bioavailability PGP Inhibition CYP2C19 Inhibition CYP2D6 Inhibition
GIN (no pretrain) 93.5 $\pm$ 1.7 67.1 $\pm$ 1.1 63.4 $\pm$ 0.9 89.1 $\pm$ 0.6 87.4 $\pm$ 0.3 84.8 $\pm$ 0.4
L2P-GNN 82.8 $\pm$ 2.0 50.3 $\pm$ 2.0 61.7 $\pm$ 1.9 76.6 $\pm$ 1.9 83.1 $\pm$ 0.2 80.2 $\pm$ 0.3
MGSSL (BFS) 96.6 $\pm$ 0.5 70.9 $\pm$ 0.9 62.1 $\pm$ 0.3 90.3 $\pm$ 0.6 88.0 $\pm$ 0.1 85.0 $\pm$ 0.1
MGSSL (DFS) 91.8 $\pm$ 1.0 65.4 $\pm$ 4.4 59.9 $\pm$ 3.6 87.9 $\pm$ 0.4 87.2 $\pm$ 0.2 83.6 $\pm$ 0.4
MoMu 91.1 $\pm$ 2.7 67.5 $\pm$ 2.0 61.6 $\pm$ 0.8 88.8 $\pm$ 0.4 86.4 $\pm$ 0.1 83.2 $\pm$ 0.4
MolCLR 94.0 $\pm$ 1.5 70.1 $\pm$ 2.9 65.0 $\pm$ 2.0 88.0 $\pm$ 0.6 87.1 $\pm$ 0.4 85.1 $\pm$ 0.5
G-Image 93.2 $\pm$ 1.0 69.4 $\pm$ 1.5 67.8 $\pm$ 7.0 90.5 $\pm$ 0.9 88.1 $\pm$ 0.2 85.2 $\pm$ 0.3
G-SMILES 95.5 $\pm$ 0.9 71.5 $\pm$ 2.6 64.1 $\pm$ 8.9 92.3 $\pm$ 2.9 88.5 $\pm$ 0.2 86.8 $\pm$ 0.1
G-1H NMR 94.9 $\pm$ 2.2 72.7 $\pm$ 2.6 72.3 $\pm$ 2.5 90.3 $\pm$ 2.4 88.3 $\pm$ 0.3 85.7 $\pm$ 0.6
G-13C NMR 96.5 $\pm$ 4.7 70.9 $\pm$ 2.9 67.5 $\pm$ 3.1 90.7 $\pm$ 1.9 88.4 $\pm$ 0.5 85.6 $\pm$ 0.8
G-GCMS 96.9 $\pm$ 1.6 67.5 $\pm$ 3.1 68.0 $\pm$ 2.6 91.3 $\pm$ 0.8 87.9 $\pm$ 0.2 84.3 $\pm$ 0.8
G-LCMS 92.2 $\pm$ 0.2 64.7 $\pm$ 2.5 67.9 $\pm$ 1.8 90.9 $\pm$ 0.5 87.7 $\pm$ 0.3 83.7 $\pm$ 0.3
CYP3A4 Inhibition CYP1A2 Inhibition CYP2C9 Inhibition CYP2C9 Substrate CYP2D6 Substrate CYP3A4 Substrate
GIN (no pretrain) 86.5 $\pm$ 0.5 92.0 $\pm$ 0.3 86.5 $\pm$ 0.4 56.6 $\pm$ 2.6 77.9 $\pm$ 2.8 56.7 $\pm$ 2.9
L2P-GNN 79.1 $\pm$ 0.3 88.1 $\pm$ 0.2 78.6 $\pm$ 0.3 59.9 $\pm$ 2.0 70.6 $\pm$ 2.2 51.6 $\pm$ 1.7
MGSSL (BFS) 86.8 $\pm$ 0.2 92.0 $\pm$ 0.1 86.4 $\pm$ 0.2 58.0 $\pm$ 1.4 78.6 $\pm$ 1.4 57.4 $\pm$ 2.0
MGSSL (DFS) 85.5 $\pm$ 0.2 91.5 $\pm$ 0.1 85.5 $\pm$ 0.1 59.5 $\pm$ 1.4 77.7 $\pm$ 2.5 60.9 $\pm$ 0.7
MoMu 85.6 $\pm$ 0.4 91.5 $\pm$ 0.1 86.4 $\pm$ 0.3 56.1 $\pm$ 2.1 79.8 $\pm$ 0.6 57.7 $\pm$ 2.6
MolCLR 86.6 $\pm$ 0.5 92.1 $\pm$ 0.3 86.5 $\pm$ 0.3 56.9 $\pm$ 2.3 79.9 $\pm$ 3.4 60.4 $\pm$ 1.4
G-Image 87.7 $\pm$ 0.3 92.5 $\pm$ 0.3 87.4 $\pm$ 0.3 61.9 $\pm$ 1.8 82.8 $\pm$ 1.1 65.2 $\pm$ 3.8
G-SMILES 88.0 $\pm$ 0.3 92.4 $\pm$ 0.1 88.3 $\pm$ 0.5 61.4 $\pm$ 2.8 76.8 $\pm$ 2.4 64.1 $\pm$ 2.9
G-1H NMR 87.8 $\pm$ 0.3 92.2 $\pm$ 0.2 87.8 $\pm$ 0.2 63.4 $\pm$ 1.3 79.1 $\pm$ 4.0 57.3 $\pm$ 2.3
G-13C NMR 87.5 $\pm$ 0.4 92.1 $\pm$ 0.3 87.6 $\pm$ 0.4 62.6 $\pm$ 1.6 76.9 $\pm$ 2.0 61.9 $\pm$ 4.6
G-GCMS 86.6 $\pm$ 0.3 91.3 $\pm$ 0.2 87.1 $\pm$ 0.4 60.2 $\pm$ 1.6 76.9 $\pm$ 1.4 62.4 $\pm$ 1.3
G-LCMS 87.1 $\pm$ 0.2 91.6 $\pm$ 0.2 87.0 $\pm$ 0.4 57.4 $\pm$ 2.2 73.7 $\pm$ 3.1 63.3 $\pm$ 1.8
TDC Datasets.
The results are shown in Table 5. Similar to the analysis on MoleculeNet datasets, the proposed ACML framework achieves the best performance on all datasets, demonstrating its superiority over other self-supervised pretraining strategies. Specifically, we found that G-SMILES consistently performs well on inhibition-related property tasks, achieving top or near-top results across six tasks: PGP Inhibition, CYP2C19 Inhibition, CYP2D6 Inhibition, CYP3A4 Inhibition, CYP1A2 Inhibition, and CYP2C9 Inhibition. The model’s prediction determines whether a drug inhibits specific enzymes, potentially reducing the enzyme’s ability to metabolize. Therefore, the strength of using SMILES modality can be attributed to its capability to represent functional groups and structural motifs, which effectively indicate how a molecule interacts with enzymes through binding. G-Image performs the best on two substrate tasks (which involve drugs metabolized by specific enzymes), with an average AUC-ROC of 82.8% in CYP2D6 Substrate and 65.2% in CYP3A4 Substrate, while G-1H NMR works the best in PAMPA Permeability (72.7%), Bioavailability (72.3%), and CYP2C9 Substrate (63.4%). G-GCMS achieves the highest average ROC-AUC of 96.9% in HIA, along with second best results in Bioavailability (68.0%) and PGP Inhibition (91.3%), however, it shows only trivial gains or slight declines compared to non-pretrained model on CYP2D6 Inhibition (-0.5%), CYP1A2 Inhibition (-0.7%), and CYP2D6 Substrate (-1.0%). These results suggest that no single chemical modality emerges as universally optimal, rather, each modality contributes unique advantages tailored to certain tasks, which cannot be fully substituted by others.
@@ 3 Discussion and Conclusion
In this work, we introduced the ACML framework, a novel approach tailored for molecular representational learning based on multi-modality contrastive learning. We demonstrated the inheritance of molecular graph-like structure allows graph representation to convey extensive information about molecules and treat the graph neural networks, i.e., the graph encoder, as a receptor to assimilate chemical semantics through multi-modal coordination, leading to effective and explainable molecular representational learning. Extensive experimental results on important chemical tasks, such as isomer discrimination, uncovering crucial chemical properties for drug discovery, and molecular properties prediction from MoleculeNet and TDC datasets, demonstrate the cross-modality transfer ability and the power of graph neural networks. In addition, our findings reveal a series of interesting questions to consider: of particular interest is how to effectively train a graph encoder to make it tailored to molecular representational learning. Instead of designing a deep and complex GNN framework and proposing various levels of self-supervised pretraining tasks, we demonstrate that the use of multi-modality contrastive learning can allow a shallow GNN (e.g. with no more than 5 layers) with light training resources to produce an expressive graph encoder with great interpretability. In this work, we focused primarily on small molecules, so the learning performance on polymers or biomacromolecules has not been fully explored. In addition, the chemical encoders employed pretrained weights from existing literature, derived from large and comprehensive databases, which may include molecules overlapping with the validation or test sets in downstream tasks. The graph encoder interacted solely with the chemical encoders through contrastive learning without direct exposure to these datasets.
Appendix A Methods
A.1 The ACML framework
@@
First, the inputs of each ACML framework consist of a molecular graph modality and one chemical modality. For each chemical modality, we make use of effective pre-trained encoders from publicly available sources, which have been demonstrated to be effective in the corresponding downstream tasks (See Table 6). As for molecular graph representation, we employ the graph neural network (GNN) as the graph encoder, where we elaborated several GNN variants in Section A.2. These inputs are then fed into their respective encoders to produce embeddings. Mathematically, let $g$ and $c$ denote a molecule’s graph representation and its chemical modality representation, their embeddings after passing to their corresponding encoders are:
@@
Table 6: Unimodal encoders of pervasive chemical modalities.
@@
A.2 Graph Encoder
Chemical molecules can be naturally represented as attributed relational graphs $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are the node set and the edge set. Here a node $v\in\mathcal{V}$ and an edge $(v,u)\in\mathcal{E}$ represent the atom $v$ and the chemical bond connecting $u$ and $v$, respectively. The corresponding atom features (atomic number, chirality tag, hybridization, etc,.) and bond features (bond types, stereotypes, etc,.) are treated as node attributes and edge attributes in the graph. Graph neural networks (GNNs) [60, 61] which operate on graphs, have been combined with deep learning [62] to learn representations on graph data [63, 64, 65]. GNNs are also demonstrated as the effective frameworks for molecular representational learning [66, 39].
Graph Convolution.
Many popular graph convolutional frameworks follow the message-passing scheme, including an iterative way of updating node representations based on the aggregations from neighboring nodes. Widely used approaches, such as GCN [67], GIN [57], GAT [68] and GraphSage [69], all follow message passing schemes. For the rest of the paper, we refer “message passing based GNNs” as “GNN”. For simplicity, we denote $\mathbf{x}_{u}^{(k)}$ as node feature vector of node $u$ in $k$-th layer, $\mathbf{e}_{v,u}$ as (optional) edge feature vector from node $v$ to node $u$, and let $\mathcal{N}_{u}$ represent the neighboring node set of the central node $u$. One message passing based convolution layer includes two steps: (1) create the message between the central node $u$ and its neighbor $v$, formalized by the function $\phi^{(k)}(\mathbf{x}_{u}^{(k)},\mathbf{x}_{v}^{(k)},\mathbf{e}_{v,u})$; (2) aggregate the messages of all neighbors of the central node $u$ using the operator $\bigoplus$, which denotes a differentiable and permutation invariant function, e.g., sum, mean or max operator. Mathematically, the latent node presentation is updated according to Equation 3 and will be fed into the next convolutional layer.